I'm 46 and trying to refresh my algebra skills.
How would I write an equation for the line going through the points (3,2) and (-2,4)
Thanks

Answer by rfer(16322)

first find the slope
m=(4-2)/(-2-3)
m=2/-5
m=-2/5
----------------
now slove for y intercept.
y=mx+b
2=-2/5*3+b
2=-6/5+b
2+6/5=b
b=16/5
b=3 1/5
---------------
y=-2/5x+3 1/5